# On-board Mission Replanning for Adaptive Cooperative Multi-Robot Systems

### v1

@@ Abstract
Cooperative autonomous robotic systems have significant potential for executing complex multi-task missions across space, air, ground, and maritime domains. But they commonly operate in remote, dynamic and hazardous environments, requiring rapid in-mission adaptation without reliance on fragile or slow communication links to centralised compute. Fast, on-board replanning algorithms are therefore needed to enhance resilience. Reinforcement Learning shows strong promise for efficiently solving mission planning tasks when formulated as Travelling Salesperson Problems (TSPs), but existing methods: 1) are unsuitable for replanning, where agents do not start at a single location; 2) do not allow cooperation between agents; 3) are unable to model tasks with variable durations; or 4) lack practical considerations for on-board deployment. Here we define the Cooperative Mission Replanning Problem as a novel variant of multiple TSP with adaptations to overcome these issues, and develop a new encoder/decoder-based model using Graph Attention Networks and Attention Models to solve it effectively and efficiently. Using a simple example of cooperative drones, we show our replanner consistently (90% of the time) maintains performance within 10% of the state-of-the-art LKH3 heuristic solver, whilst running 85-370 times faster on a Raspberry Pi. This work paves the way for increased resilience in autonomous multi-agent systems.
@@ I Introduction
Teams of autonomous robots, including uncrewed aerial, underwater, surface, or ground vehicles (UAVs, UUVs, USVs, UGVs) and multi-satellite constellations, have huge potential for efficiently and safely completing a diverse range of complex missions. In comparison to single-agent operations, such multi-agent systems have increased robustness to failure [1], greater areal coverage [2], and faster mission completion [3]. Furthermore, they can cooperatively tackle complex missions comprised of multiple tasks, where these tasks can be as diverse as capturing images from space [4], manufacturing or repairing items or transporting materials [5], or making environmental measurements [6]. However, these systems also commonly operate in environments that are both remote or isolated and also highly dynamic and hazardous, e.g. inside nuclear power plants [7], over active volcanoes [6] or in congested low earth orbits in space [4]. Operating in such challenging environments requires these systems to be able to adapt and replan in response to major unexpected changes that impact the mission plan [8] such as the identification of new tasks or the loss or failure of a robotic agent, without relying on fragile or slow communication links to centralized ground stations [9] (Figure 1a). Therefore, lightweight and fast replanning algorithms that can be deployed on-board robotic systems are needed to enable autonomous adaptability during a mission.
@@
[13] Capacitated VRP Graph Attention Attention Model REINFORCE (Rollout) $\times$ $\times$ $\times$ $\times$ $\times$
[14] multi-tasks VRP Graph Attention Attention Model POMO + MoE Layer $\times$ $\times$ $\times$ (Goal) $\times$
@@
[16] mTSP with Multiple Depots MoE Net. Ptr. Net. Pretrained $\times$ $\times$ (Agent)
@@
(Task, Agent, Collab.)
*VRP refers to the Vehicle Routing Problem; Ptr. Net. refers to the Pointer Network; MoE refers to the Mixture of Experts technique;
@@
Because mission planning with multiple agents requires finding the optimal tour path for each agent while minimizing the length of the longest tour, previous studies have often formulated and solved it as a variant of the multiple Traveling Salesman Problem (mTSP) with a min-max objective [18]. However, our particular planning problem has three major differences to traditional mTSP (Figure 1b), and in addition requires practical consideration for edge deployment onboard robotic systems. First, variable start locations for agents must be considered, as each agent may be at any location within the mission domain when replanning is required. Second, nodes with variable time costs are necessary to reflect the common real-world situation where mission time depends not just on how long it takes for an agent to travel to a task, but also how long it takes to complete that task. This models a wide range of applications where tasks have differing execution times [19, 20, 21]. Third, cooperative tasking is required, such that multiple agents can work together on a single task to complete it faster, even if this joint contribution is asynchronous rather than simultaneous. No previous studies have addressed this important combination of adaptations (see “Key Adaptation” columns of Table I), which we name the Cooperative Mission Replanning Problem (CMRP). Finally, to enable onboard replanning, any model for solving CMRP must be suitably small and fast [22]. These limitations also mean that a single generalised model that can solve a range of problem sizes is needed, as the number of available agents and remaining tasks are likely to vary during a mission [23] and it is impractical to store one model per problem size on memory-constrained edge devices.
@@
The mTSP class of problems are NP-hard [24] with the total number of possible mission plans scaling as $\frac{\left(n+m-1\right)!}{2\left(m-1\right)!}$, where $n$ is the number of tasks and $m$ is the number of agents. Reinforcement Learning (RL) techniques, especially those using Attention mechanisms (e.g. GAT; Graph Attention Networks), have proven to be effective at overcoming these scaling issues and efficiently solving TSP [10, 11, 12], mTSP [25], or other similar multi-agent problems [13, 14]. These methods are able to scale to larger problems than exact solvers such as CPLEX [26], are faster at execution time and generalise better than heuristic methods [27, 25, 28] such as the Lin-Kernighan heuristic [29], show more consistent and generalisable performance than metaheuristic methods [30, 31], and can be deployed on edge devices as lightweight models [32]. However, no work to date has addressed the replanning problem as outlined above. Several studies have made one of the required adaptations, e.g. multiple depots, where agents start and finish at different locations  [15, 16] or cooperative tasking [17], but none have considered more than one together, and no previous studies have considered nodes with variable time costs (Table I), which is a fundamental requirement. Similarly, most prior research has focused on high-performance computing systems with GPUs, with limited investigation of deployment on ground robots [16] or CPUs [17]. And whilst [14, 17] developed a single neural solver for different problem variants or numbers of agents, they did not consider varying numbers of tasks. Our work aims to fill this critical gap, developing a cooperative replanning solver that is suitable for edge deployment and real-world on-board application.
@@
  • To our knowledge, we are the first to formulate the Cooperative Mission Replanning Problem (CMRP); a novel variant of the mTSP adapted to include variable agent start locations, nodes with variable time cost, and cooperative tasking.
@@
  • We propose a lightweight mission replanner based on the GAT and Attention Model that can effectively and efficiently solve the CMRP.
@@
  • Using an example of aerial drones completing a set of tasks in a 2D domain, our approach produces solutions that are consistently within 10% of those generated by the state-of-the-art LKH3 heuristic solver, but which runs 85-370 times faster than the same solver on a Raspberry Pi. We also demonstrate that a single generalised model performs well across a range of problem sizes, with minimal degradation.
@@
Overall, our new formulation and approach show the potential to increase the resilience of a wide variety of multi-agent robotic systems, through enabling fast and robust in-mission, on-device replanning.
@@
Figure 2: (a) The RL training process. (b) The transformation of the mission planning data into a higher-level node embeddings and a graph embedding using a GAT encoder. (c) The Attention Model decoder generate a probability distribution for avalibale agent-task combinations based on the graph embedding, the current node embedding, the current state of the mission (in embedding space) and the avalibility mask. The agent-task step with the highest probability is then selected as the next step. (d) The sequential generation of all steps of the mission plan.
@@ II-A The Cooperative Mission Replanning Problem (CMRP)
Traditional mTSP with $m$ salespeople or agents and $n$ cities or tasks can be defined as a graph $G=(V,E)$ where the vertices $V=\{v_{depot},v_{1},...v_{n}\}$ include all $n$ task locations as well as $v_{depot}$, which serves as the single start and end location for all agents. The edge set $E$ is defined as $E=\{(v_{i},v_{j}):v_{i},v_{j}\in V,i\neq j\}$ and represents all possible valid paths between tasks. For all edges, $T=(d_{ij})$ represents the travel time associated with transversing from task $v_{i}$ to $v_{j}$, assuming a constant travel speed of $1$ m/s.
@@
Second, we incorporate variable time costs for each task so that each vertex $v_{i}$ has an associated non-negative time cost $t_{i}$, with the exception for the depot and the start locations, where $t_{i}=0\text{ for }v_{i}\in\{v_{depot},v_{start1},...v_{startm}\}$. This is similar to the nodes’ demand in Capacitated VRP [13] with the distinction that $t_{i}$ is incorporated into the edge costs rather than being a separated vector. Our new cost matrix is therefore calculated as $T^{\prime}=d_{ij}+t_{j}$, and again this modification introduces more asymmetry to this matrix, as time costs are no longer determined just by the distance between two vertices, but also by the direction of travel.
@@
These combined changes from $V$, $T$, $n$, $t$ to $V^{\prime}$, $T^{\prime}$, $n^{\prime}$, $t^{\prime}$ respectively define our problem as a variant of mTSP with $m$ agents and $n^{\prime}+m+1$ vertices (including the home depot; see Figure 1). Hence, the number of unique solutions (i.e. possible mission plans) is:
@@ II-B The Graph Attention Replanner (GATR)
Our proposed GAT-based mission replanner (GATR) builds upon the encoder-decoder architecture from [13]. We adapt the mTSP environment from the RL4CO Library [25] to incorporate different agents’ start locations and time costs associated with nodes (the first and second changes to standard mTSP described in Table I). Discretization of tasks to allow splitting of tasks between cooperative agents (the third change) does not require any additional adaptation, and can simply be accommodated during problem setup by replacing a single task of a certain time cost with several shorter collocated tasks. Finally, we enhance the generator and data ingestion pipeline to enable training of a single generalised model for multiple problem sizes.
The overall training process is shown in Figure 2a. Using the environment described above, we trained our model using the simple REINFORCE algorithm with a greedy rollout baseline [13] as detailed in Appendix -B. There is a possibility to extend this to other algorithms, which we leave for future work. The reward is defined as the negative of the maximum (largest) mission time among the agents, as in [14].
@@ II-B1 Encoder
The primary role of the GAT encoder is to transform an input graph into latent representations, as illustrated in the first step of the flowchart in Figure 2a. In addition to the standard mTSP graph features (depot locations and task locations), we also include the time cost of vertices and agents’ starting locations as part of the input to the encoder. This is similar to the approach used for specifying the ‘demand’ for deliveries in the Capacitated VRP [13], and for incorporating multiple depots in [15]. The GAT encoder then transform these features into a set of node embeddings and a graph embedding that are passed to the decoder using the procedure described in Appendix -A2.
@@ II-B2 Decoder
Our decoder is based on the Attention Model, which was selected over the traditional Pointer Network due to its empirically superior performance [12]. The decoder forms the mission plan sequentially, as illustrated in Figure 2c-d. At each time step, the decoder selects the next sub-task to visit for an agent based on the graph embedding, the current node embedding, the current state of the mission (in embedding space) and the availability mask. If the depot is selected, the algorithm proceeds to the subsequent agent. Note that an agent has the option to stop completing tasks and return to the depot at any point, provided there is still at least one other agent available to complete remaining tasks, similar to the assumption made in [16]. The state of the environment is summarized by the following features: the number of remaining agents, the accumulated mission time for the current agent, the maximum mission time across all agents, and the distance from the depot. These features are mapped into an embedding space using affine linear transformations and concatenated with the embedding for the current agent vertex (i.e. the agent’s current location) to form the context embeddings. Unlike in [13], we do not include the embedding of the first vertex traversed by the agent, as the agent does not return to this location in the replanning problem. Cross-attention with an availability mask is applied to the context embeddings and vertex embeddings to compute the attention scores. In the replanning problem, the mask prevents selection of:
  • Previously visited sub-tasks
  • Start locations when the current node is not the depot
  • The depot if the current agent is the last available and unvisited sub-tasks remain
The next sub-task is then selected based on these scores, as detailed in Appendix -A3. This process is repeated until all sub-tasks are assigned to the agents. The final output is a set of routes, one for each drone, covering all sub-tasks, as illustrated in Figure 2d.
@@ III Results
We first demonstrate GATR’s effectiveness for the CMRP on small problem sizes where it is possible to evaluate our replan against all possible solutions (Section III-A). Then we show how a single generalised GATR model is able to solve problems with varying numbers of tasks, agents, and task discretization levels, and how our model scales to larger problem sizes (Section III-B).
@@ III-A Analysis for small problem sizes
Figure 3: Comparison of how GATR and other benchmarks (columns) solve an example three-agent-four-task problem for our five progressively complex problem types (rows). The tours of the three agents are shown in red, amber, and green, the tasks are shown in black circles (with size corresponding to the time cost of each task for the last three rows), and the black square shows the home depot. In second, fourth and fifth rows, where agent start locations differ from the depot, they are marked with a black triangle.
@@
Figure 4: (a) Histograms showing the distribution of mission times for our method GATR (top row) and four benchmarks for the test set of 300 scenarios for the three-agent-four-task CMRP. The circled letters represent the five different solutions for the single example scenario shown in the bottom row of Fig. 3. (b) A cumulative frequency histogram showing the percentage of GATR solutions (y-axis) with normalized mission times less than a given value (x-axis). The solid lines represent comparison against the Opt-P benchmark ($\hat{mt}_{Opt-P}$) for the five progressively complex problem types illustrated in Fig. 3, building up from mTSP (pink) to CMRP, our full problem (dark blue). The dashed dark blue line represents comparison against NOpt-LKH3 benchmark ($\hat{mt}_{NOpt-LKH3}$) for CMRP. (c) A graph illustrating the proportion of solutions with $\hat{mt}$ below the 1% threshold for different problem types and numbers of sub-tasks. The triangles represent the four-task problems and match the colours for the same set of experiments shown in (b).
@@
Our experimental setup simulates a small indoor flight arena for aerial drones, with a 10 m×$\times$10 m 2D domain. Task, depot and agent start locations are randomly distributed, and task time costs are randomly set between 1-10s. We assume agents move at a fixed speed of 1 m/s. In order to systematically assess GATR’s capabilities, we incrementally incorporate the changes to mTSP from Section II-A, defining five problem types: mTSP; mTSP with multiple agent start locations; mTSP with varying task time costs; mTSP with both multiple start locations and varying task time costs; and finally our CMRP problem which also features agent cooperation (tasks discretised into multiple sub-tasks). Initially, we focus on a three agent, four task problem, and train five models (one for each problem type), using three different seeds for each. We use task discretization level $\delta=2$ for collaborative tasking, which means that our CMRP has 8 sub-tasks (4 tasks ×$\times 2$) whilst all other problem types have 4 sub-tasks (only 1 sub-task per task). We generate 100 mTSP evaluation instances also across three seeds (300 instances in total). We use mission time (time to complete all tasks and return to homebase) as our primary evaluation metric.
@@
We evaluate mission times against four benchmarks: the optimal and median population solutions (Opt-P and Med-P respectively), calculated via brute-force enumeration of the full population of possible solutions (e.g. 6,652,800 for 3 agents and 8 sub-tasks, from Equation 1); the median of a sample of the population - 101 randomly-generated solutions (Med-S); and a near-optimal solution generated by the LKH3 solver [29] (NOpt-LKH3). Opt-P and NOpt-LKH3 serve as the upper bounds on expected performance, whilst the two median benchmarks represent a lower bound as they simulate the expected performance of a randomly generated plan. NOpt-LKH3 and Med-S are essential for evaluation of large problem sizes where enumeration of all solutions is not possible.
@@
Qualitative inspection of example solutions (e.g. single example scenario in Fig. 3) shows that GATR exhibits sensible and desired behaviours across all five problem types: effectively distributing tasks between agents, selecting tasks close to agent locations, accounting for variable task time cost when assigning agents, and collaboration on tasks.
@@
This is supported by analysis of the full test set of 300 scenarios; GATR solutions in general have mission times similiar to those from Opt-P and NOpt-LKH3 solutions (Fig. 4a) and much smaller than those from Med-P and Med-S solutions. In addition, the close alignment between Med-S and Med-P distributions in Fig. 4a provides confidence in approximating the median of the full population with a sample-based estimate for evaluation of larger problem sizes.
@@
We also analyse normalised mission time ($\hat{mt}_{benchmark}$), where GATR’s mission times ($mt_{GATR}$) are compared against those from benchmark solutions:
@@
where $mt_{benchmark}$ can refer to either mission time from the optimal benchmark ($mt_{Opt-P}$) or the near-optimal benchmark ($mt_{NOpt-LKH3}$) as desired. Here, $\hat{mt}_{benchmark}=0$ indicates our model produces a solution with the same mission time as the chosen benchmark (the optimal solution or a near-optimal solution from the LKH3 solver); and $\hat{mt}_{benchmark}=1$ indicates our model’s solution has the same mission time as an average randomly generated mission plan. Fig. 4b shows that 89.3% of GATR solutions achieve $\hat{mt}_{Opt-P}<0.1$ (normalised mission times within 10% of optimal), indicating consistently high performance despite the increased problem complexity over mTSP, and 12% solutions actually outperform our LKH3 heuristic benchmark (evidenced by the dashed line to the left of 0.00 in Fig. 4b).

### v2

# Onboard Mission Replanning for Adaptive Cooperative Multi-Robot Systems

@@ Abstract
Cooperative autonomous robotic systems have significant potential for executing complex multi-task missions across space, air, ground, and maritime domains. But they commonly operate in remote, dynamic and hazardous environments, requiring rapid in-mission adaptation without relying on fragile or slow communication links to centralized compute. Fast, on-board replanning algorithms are therefore essential to enhance resilience for these systems, but do not yet exist. Reinforcement Learning (RL) shows strong promise for efficiently solving mission planning tasks formulated as Travelling Salesperson Problems (TSPs), but existing methods: 1) are unsuitable for replanning, where agents do not start at a single location; 2) do not allow cooperation between agents; 3) are unable to model tasks with variable durations; or 4) lack practical considerations for on-board deployment. Here we address this gap by defining the Cooperative Mission Replanning Problem as a novel adaptation of multiple TSP, and develop a new encoder/decoder-based RL model to solve it effectively and efficiently. Using a simple example of cooperative drones, we show our replanner consistently (90% of the time) maintains performance within 10% of the state-of-the-art LKH3 heuristic solver, whilst running 85-370 times faster on a Raspberry Pi. This work paves the way for increased resilience in autonomous multi-agent systems.
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: TBC
@@ I Introduction
Teams of autonomous robots, including uncrewed aerial, underwater, surface, or ground vehicles (UAVs, UUVs, USVs, UGVs) and multi-satellite constellations, have huge potential for efficiently and safely completing a diverse range of complex missions. In comparison to single-agent operations, such multi-agent systems have increased robustness to failure [1], greater areal coverage [2], and faster mission completion [3]. Furthermore, they can cooperatively tackle complex missions comprised of multiple tasks, where these tasks can be as diverse as capturing images [4], manufacturing/ repairing items or transporting materials [5], or making environmental measurements [6]. However, these systems also commonly operate in environments that are both remote or isolated and also highly dynamic and hazardous, e.g. in nuclear power plants [7], over active volcanoes [6] or in congested low earth orbits in space [4]. Operating in such challenging environments requires these systems to be able to adapt and replan in response to major unexpected changes that impact the mission plan [8] such as the identification of new tasks or the loss or failure of a robotic agent, without relying on fragile or slow communication links to centralized ground stations [9] (Fig. 1a). Therefore, lightweight and fast replanning algorithms that can be deployed on-board cooperative robotic systems are critical for enabling autonomous adaptability during a mission; but these do not yet exist.
@@
[13] CVRP Graph Attention Attention Model REINFORCE (Rollout) $\times$ $\times$ $\times$ $\times$ $\times$
[14] multi-tasks VRP Graph Attention Attention Model POMO + MoE Layer $\times$ $\times$ $\times$ ✓(Goal) $\times$
@@
[16] mTSP with Multiple Depots MoE Net. Ptr. Net. Pretrained $\times$ $\times$ ✓(Agent)
@@
[18] VRP (Min-Max) Graph Attention (Partition/Navigation) Attention Model REINFORCE (Rollout) $\times$ $\times$ ✓(Depot) $\times$
[19] Heterogeneous CVRP (Min-Max) Graph Attention Attention Model (Parallel) REINFORCE $\times$ $\times$ ✓(Agent) $\times$
@@
✓(Task, Agent, Collab.)
VRP refers to the Vehicle Routing Problem; CVRP refers to the Capacitated VRP; Ptr. Net. refers to the Pointer Network; MoE refers to the Mixture of Experts technique;
@@
Because mission planning with multiple agents requires finding the optimal tour path for each agent while minimizing the length of the longest tour, previous studies have often formulated and solved it as a variant of the multiple Traveling Salesman Problem (mTSP) with a min-max objective [20]. However, our particular planning problem has three major differences to traditional mTSP (Fig. 1b), and in addition requires practical consideration for edge deployment onboard robotic systems. First, variable start locations for agents must be considered, as each agent may be at any location within the mission domain when replanning is required. Second, nodes with variable time costs are necessary to reflect the common real-world situation where mission time depends not just on how long it takes for an agent to travel to a task, but also how long it takes to complete that task. This models a wide range of applications where tasks have differing execution times [21, 22, 23]. Third, cooperative tasking is required, such that multiple agents can work together on a single task to complete it faster, even if this joint contribution is asynchronous rather than simultaneous. No previous studies have addressed this important combination of multiple adaptations (see “Key Adaptation” columns of Table I), which we name the Cooperative Mission Replanning Problem (CMRP). Finally, to enable onboard replanning, any model for solving CMRP must be suitably small and fast [24]. These limitations also necessitate a single generalized model that can solve a range of problem sizes, as the number of available agents and remaining tasks are likely to vary during a mission [25] and it is impractical to store one model per problem size on edge devices.
The mTSP class of problems are NP-hard [26] with the total number of possible mission plans scaling as $\frac{\left(n+m-1\right)!}{2\left(m-1\right)!}$, where $n$ is the number of tasks and $m$ is the number of agents. Reinforcement Learning (RL) techniques, especially those using Attention mechanisms (e.g. GAT; Graph Attention Networks), have proven to be effective at overcoming these scaling issues and efficiently solving TSP [10, 11, 12], mTSP [27], min-max Vehicle Routing Problems (VRP) [18, 19], or other similar multi-agent problems [13, 14].
These methods are able to scale to larger problems than exact solvers such as CPLEX [28], are faster at execution time and generalize better than heuristic methods [29, 27, 30] such as the Lin-Kernighan heuristic [31], show more consistent and generalizable performance than metaheuristic methods [32, 33], and can be deployed on edge devices as lightweight models [34].
@@
But despite its real-world importance, no work to date has addressed the replanning problem as outlined above. Several studies have made one of the required adaptations, e.g. multiple depots, where agents start and finish at different locations [15, 16, 18] or cooperative tasking [17], but none have considered more than one together, and no previous studies have considered nodes with variable time costs (Table I), which is a fundamental requirement. Similarly, most prior research has focused on high-performance computing systems with GPUs, with limited investigation of deployment on ground robots [16] or CPUs [17]. And whilst [14, 17, 18, 19], developed a single neural solver for different problem variants, numbers of agents or depots, they did not consider varying numbers of tasks. Our work fills this critical gap, developing a cooperative replanning solver that is suitable for edge deployment and real-world on-board application.
@@
  • We are the first to formulate the Cooperative Mission Replanning Problem (CMRP); a novel and important real-world variant of the mTSP adapted to include variable agent start locations, nodes with variable time cost, and cooperative tasking.
@@
  • We propose a new lightweight mission replanner based on the GAT and Attention Model that can effectively and efficiently solve the CMRP on edge hardware.
@@
  • Using an example of UAVs completing tasks in a 2D domain, our approach produces solutions that are consistently within 10% of those generated by the state-of-the-art LKH3 heuristic solver, but runs 85-370 times faster on a Raspberry Pi. Our single generalized model has high peformance across a range of problem sizes.
  • Our framework also valuably supports future extension to additional important capabilities (e.g. multi-objective, probabilistic, anticipatory replanning).
@@
Our new formulation and approach enables fast and robust in-mission, on-device replanning for the first time, for cooperative teams of robots. This has the potential to improve the resilience of diverse real-world multi-agent robotic systems.
@@
Figure 2: (a) The RL training process. (b) The transformation of the mission planning data into higher-level node embeddings and a graph embedding using a GAT encoder. (c) The Attention Model decoder generates a probability distribution for available agent-task combinations based on the graph embedding, the current node embedding, the current state of the mission (in embedding space) and the availability mask. The agent-task step with the highest probability is then selected as the next step. (d) The sequential generation of all steps of the mission plan.
@@ II-A The Cooperative Mission Replanning Problem (CMRP)
Traditional mTSP with $m$ salespeople or agents and $n$ cities or tasks can be defined as a graph $G=(V,E)$ where the vertices $V=\{v_{depot},v_{1},...v_{n}\}$ include all $n$ task locations as well as $v_{depot}$, which serves as the single start and end location for all agents. The edge set $E$ is defined as $E=\{(v_{i},v_{j}):v_{i},v_{j}\in V,i\neq j\}$ and represents all possible valid paths between tasks. For all edges, $T=(d_{ij})$ represents the travel time associated with traversing from task $v_{i}$ to $v_{j}$, assuming a constant travel speed of $1$ m/s.
@@
Second, we incorporate variable time costs for each task so that each vertex $v_{i}$ has an associated non-negative time cost $t_{i}$, with the exception of the depot and the start locations, where $t_{i}=0\text{ for }v_{i}\in\{v_{depot},v_{start1},...v_{startm}\}$. This is similar to the nodes’ demand in Capacitated VRP (CVRP) [13] with the distinction that $t_{i}$ is incorporated into the edge costs rather than being a separate vector. Our new cost matrix is therefore calculated as $T^{\prime}=d_{ij}+t_{j}$, and again this modification introduces more asymmetry to this matrix, as time costs are no longer determined just by the distance between two vertices, but also by the direction of travel.
@@
These combined changes from $V$, $T$, $n$, $t$ to $V^{\prime}$, $T^{\prime}$, $n^{\prime}$, $t^{\prime}$ respectively define our problem as a variant of mTSP with $m$ agents and $n^{\prime}+m+1$ vertices (including the home depot; see Fig. 1). Hence, the number of unique solutions (i.e. possible mission plans) is:
@@ II-B The Graph Attention Replanner (GATR)
Our proposed GAT-based mission replanner (GATR) builds upon the encoder-decoder architecture from [13]. We adapt the mTSP environment from the RL4CO Library [27] to incorporate different agents’ start locations and time costs associated with nodes (the first and second changes to standard mTSP described in Table I). Discretization of tasks to allow splitting of tasks between cooperative agents (the third change) does not require any additional adaptation, and can simply be accommodated during problem setup by replacing a single task of a certain time cost with several shorter collocated tasks. Finally, we enhance the data ingestion pipeline to enable training of a single generalized model for multiple problem sizes.
The overall training process is shown in Fig. 2a. Using the environment described above, we trained our model using the REINFORCE algorithm with a greedy rollout baseline [13] (see Appendix). The reward is defined as the negative of the maximum (largest) mission time among the agents as in [18].
@@ II-B1 Encoder
The primary role of the GAT encoder is to transform an input graph into latent representations, as illustrated in the first step of the flowchart in Fig. 2a. In addition to the standard mTSP graph features (depot locations and task locations), we also include the time cost of vertices and agents’ starting locations as part of the input to the encoder. This is similar to the approach used for specifying the ‘demand’ for deliveries in the CVRP [13], and for incorporating multiple depots in [15]. The GAT encoder then transforms these features into a set of node embeddings and a graph embedding that are passed to the decoder using the procedure described in the Appendix.
@@ II-B2 Decoder
Our decoder is based on the Attention Model, which was selected over the traditional Pointer Network due to its empirically superior performance [12]. The decoder forms the mission plan sequentially, as illustrated in Fig. 2c-d. At each time step, the decoder selects the next sub-task to visit for an agent based on the graph embedding, the current node embedding, the current state of the mission (in embedding space) and the availability mask. If the depot is selected, the algorithm proceeds to the subsequent agent. Note that an agent has the option to stop completing tasks and return to the depot at any point, provided there is still at least one other agent available to complete remaining tasks, similar to the assumption made in [16]. The state of the environment is summarized by the following features: the number of remaining agents, the accumulated mission time for the current agent, the maximum mission time across all agents, and the distance from the depot. These features are mapped into an embedding space using affine linear transformations and concatenated with the embedding for the current agent vertex (i.e. the agent’s current location) to form the context embeddings. Unlike in [13], we do not include the embedding of the first vertex traversed by the agent, as the agent does not return to this location in the replanning problem. Cross-attention with an availability mask is applied to the context embeddings and vertex embeddings to compute the attention scores. In the replanning problem, the mask prevents selection of: previously visited sub-tasks; start locations when the current node is not the depot; and the depot if the current agent is the last one available, and unvisited sub-tasks remain. The next sub-task is then selected based on these scores, as detailed in the Appendix. This process is repeated until all sub-tasks are assigned to the agents. The final output is a set of routes, one for each agent, covering all sub-tasks, as illustrated in Fig. 2d. We note that in future work, the GATR architecture could be easily extended, e.g. to other encoders tailored for min-max objectives (e.g. [18], which has separate embeddings for task partitioning and navigation), or decoders that output actions for agents in parallel to encourage collaborative behavior [19].
@@ III Results
We first demonstrate GATR’s effectiveness for the CMRP on small problem sizes where it is possible to evaluate our replan against all possible solutions (Section III-A). Then we show how a single generalized GATR model is able to solve problems with varying numbers of tasks, agents, and task discretization levels, and how our model scales to larger problem sizes (Section III-B).
@@ III-A Analysis for small problem sizes
Figure 3: Comparison of how GATR and other benchmarks (columns) solve an example three-agent-four-task problem for our five increasingly complex problem types (rows). The tours of the three agents are shown in red, amber, and green, the tasks are shown in black circles (with size corresponding to the time cost of each task for the last three rows), and the black square shows the home depot. In second, fourth and fifth rows, where agent start locations differ from the depot, they are marked with a black triangle.
@@
Figure 4: (a) Histograms showing the distribution of mission times for our method GATR (top row) and four benchmarks for the test set of 300 scenarios for the three-agent-four-task CMRP. The circled letters represent the five different solutions for the single example scenario shown in the bottom row of Fig. 3. (b) A cumulative frequency histogram showing the percentage of GATR solutions (y-axis) with normalized mission times less than a given value (x-axis). The solid lines represent comparison against the Opt-P benchmark ($\hat{mt}_{Opt-P}$) for the five progressively complex problem types illustrated in Fig. 3, building up from mTSP (pink) to CMRP, our full problem (dark blue). The dashed dark blue line represents comparison against NOpt-LKH3 benchmark ($\hat{mt}_{NOpt-LKH3}$) for CMRP. (c) A graph illustrating the proportion of solutions with $\hat{mt}$ below the 0.01 (1%) threshold for different problem types and numbers of sub-tasks. The triangles represent the four-task problems and match the colors for the same set of experiments shown in (b).
@@
Our experimental setup simulates a small indoor flight arena for aerial drones, with a $\times10 m$ 2D domain. Task, depot and agent start locations are randomly distributed, and task time costs are randomly set between 1-10 s. We assume agents move at a fixed speed of 1 m/s. In order to systematically assess GATR’s capabilities, we incrementally incorporate the changes to mTSP from Section II-A, defining five problem types: mTSP; mTSP with multiple agent start locations; mTSP with varying task time costs; mTSP with both multiple start locations and varying task time costs; and finally our CMRP problem which also features agent cooperation (tasks discretized into multiple sub-tasks). Initially, we focus on a three agent, four task problem, and train five models (one for each problem type), using three different seeds for each. We use task discretization level $\delta=2$ for collaborative tasking, which means that our CMRP has 8 sub-tasks (4 tasks $\times 2$) whilst all other problem types have 4 sub-tasks (only 1 sub-task per task). We generate 100 mTSP evaluation instances also across three seeds (300 instances). We use mission time (time to complete all tasks and return to homebase) as our primary evaluation metric.
@@
We evaluate mission times against four benchmarks: the optimal and median population solutions (Opt-P and Med-P respectively), calculated via brute-force enumeration of the full population of possible solutions (e.g., 1,814,400 for 3 agents and 8 sub-tasks, from Equation 1, evaluation of which takes $\sim$1 minute for a single scenario on the GPU); the median of a sample of the population - 101 randomly-generated solutions (Med-S); and a near-optimal solution generated by the LKH3 solver [31] (NOpt-LKH3). Opt-P and NOpt-LKH3 serve as the upper bounds on expected performance, whilst the two median benchmarks represent a lower bound as they simulate the expected performance of a randomly generated plan. We note that although it is feasible to obtain enumeration solutions for these small problem sizes, this is not possible for the larger problems we investigate in Section III-B, for which NOpt-LKH3 and Med-S are essential for evaluation instead.
@@
Qualitative inspection of example solutions (e.g. single example scenario in Fig. 3) shows that GATR exhibits sensible and desired behaviors across all five problem types: effectively distributing tasks between agents, selecting tasks close to agent locations, accounting for variable task time cost when assigning agents, and collaboration on tasks.
@@
This is supported by analysis of the full test set of 300 scenarios; GATR solutions in general have mission times similar to those from Opt-P and NOpt-LKH3 solutions (Fig. 4a) and much smaller than those from Med-P and Med-S solutions. In addition, the close alignment between Med-S and Med-P distributions in Fig. 4a provides confidence in approximating the median of the full population with a sample-based estimate for evaluation of larger problem sizes.
@@
We also analyse normalized mission time ($\hat{mt}_{benchmark}$), where GATR’s mission times ($mt_{GATR}$) are compared against those from benchmark solutions: